### 10 Jan 2024 v1

# INACIA: Integrating Large Language Models in Brazilian Audit Courts: Opportunities and Challenges

Abstract.
This paper introduces INACIA (Instrução Assistida com Inteligência Artificial), a groundbreaking system designed to integrate Large Language Models (LLMs) into the operational framework of Brazilian Federal Court of Accounts (TCU). The system automates various stages of case analysis, including basic information extraction, admissibility examination, Periculum in mora and Fumus boni iuris analyses, and recommendations generation. Through a series of experiments, we demonstrate INACIA’s potential in extracting relevant information from case documents, evaluating its legal plausibility, and generating judicial recommendations. Utilizing a validation dataset alongside LLMs, our evaluation methodology presents an innovative approach to assessing system performance, correlating highly with human judgment. The results highlight INACIA’s proficiency in handling complex legal tasks, indicating its suitability for augmenting efficiency and judicial fairness within legal systems. The paper also discusses potential enhancements and future applications, positioning INACIA as a model for worldwide AI integration in legal domains.
copyright: nonejournalyear: 2023doi: XXXXXXX.XXXXXXXjournal: DGOVjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8
1. Introduction
Large Language Models (LLMs) have emerged as a pivotal innovation in the artificial intelligence space, altering our interaction with digital technologies. Their applications span a diverse spectrum, including but not limited to information retrieval, data extraction, text classification, machine translation, and question answering (Kalyan, 2023). Beyond these conventional domains, LLMs demonstrate remarkable capabilities in areas such as standardized testing, creative writing (Shanahan and Clarke, 2023), and even complex decision-making tasks in business and legal environments (Sun, 2023; Chen et al., 2023). This versatility showcases their ability to generate insights and solutions that are contextually relevant.
In this paper, we introduce a groundbreaking application of LLMs for administrative case processing and autonomous decision-making developed jointly by the Brazilian Federal Court of Accounts (TCU, or Tribunal de Contas da União, in Portuguese) and experts at several Brazilian research organizations. INACIA (Instrução Assistida com Inteligência Artificial, in Portuguese or Instruction Assisted with Artificial Intelligence) is a system that analyzes and issues decisions [^1] These decisions are somewhat similar to expert opinions but rather than being offered by external experts, these are part of the responsibilities of staff auditors at TCU. They are “decisions” in the sense that they can be adopted in full and are binding once staff auditors and division heads sign off on the recommended course of action. for a subset of cases under TCU’s responsibility. Its inputs are case filings; outputs are (official) court decision documents and tabular data required for the court’s record-keeping activities. INACIA uses a combination of LLMs and search algorithms to extract information from case documents and to reason through the extracted information.
The main challenge in developing systems like INACIA is evaluating their performance. How can we ensure the quality of the produced decisions and extracted data? To address this challenge, we construct a validation dataset from TCU’s case filings and propose a methodology that uses LLMs to evaluate the system’s outputs. To our knowledge, this is the first evaluation of this kind in the legal domain. We demonstrate that our methodology is highly correlated with human evaluation and can be used to evaluate the system’s performance. Our results demonstrate that INACIA can produce high-quality decisions but also forewarn the need for further research to improve the system’s performance.
2. Background on Brazilian Audit Courts
The Brazilian Federal Court of Accounts is the external control institution that supports the National Congress with the mission of overseeing the budget and the financial execution of the Federal Government. TCU issues rulings that serve to prevent deviations from established norms, offers guidance, recommends enhancements, and, when deemed necessary, imposes administrative sanctions on government agencies and agents implicated in any form of misconduct.
2.1. Role and Responsibilities
In most contemporary democracies, a network of institutions exists to regulate, investigate, and potentially sanction individuals occupying elected and non-elected office for improperly using public funds (Speck, 2011). Despite corruption often dominating media attention, mismanagement of public resources makes up the majority of audit findings across Brazil (Ferraz and Finan, 2008). TCU occupies a premier position within the control network of the Brazilian government and is seen as a model institution for auditing practices across the country. [^2] Each state in Brazil also has its state audit court heavily inspired by TCU and closely tied to the Federal Court.
2.2. Artificial Intelligence and TCU
Artificial Intelligence (AI) technologies are not new to TCU. Over the past five years, court developers have been working on systems to summarize case documents, to efficiently search and retrieve case law information to support decisions, and, more recently, to create an internal ChatGPT-like conversational bot to help staff find all kinds of unstructured, internal information. Earlier in 2023, TCU created its own Artificial Intelligence Center (or Núcleo de Inteligência Artificial, in Portuguese), whose goal is to guide and implement responsible, ethical use of AI in the court.
Recognizing the rapid evolution of AI systems, large language models in particular, TCU issued a call for proposals in early 2022, inviting external vendors to submit proposals to develop an end-to-end system to streamline audit case processing. TCU aims to focus auditors’ expertise on the most challenging tasks of case processing while handing over the more routine tasks to the AI system. This strategy is well-aligned with the evidence on how best to incorporate technology for productivity gains and labor output augmentation (Acemoglu and Autor, 2011).
2.3. Audit Case Processing
TCU investigates (and eventually issues rulings) the use of federal funds in two ways: (1) when TCU independently audits federal agencies (or anyone using federal funds) or (2) when TCU receives complaints filed by individuals, private entities, or public agencies on the misuse of federal funds.
While case proceedings may vary based on the initiation method, case processing is typically a time-intensive task primarily handled by TCU staff. Auditors have to review case filings, confirm whether TCU should hear the case, search and append external evidence of wrongdoing (if necessary), check claims against case law and federal statutes, apply the Court’s jurisprudence and prepare a final document summarizing everything, including the recommended action on the case, to be sent over to a panel of judges for final adjudication. TCU estimates that reviewing a single case requires approximately 30 hours, incurring a cost of approximately \$1,750 per case before a case is even presented to the court’s judicial panel. Given an annual average processing rate of about 2,000 cases and the textual nature of the task, audit case processing is an ideal candidate for an LLM application that would help expedite review while reducing costs for the organization. [^3] For comparison purposes, INACIA processes a case in under 20 minutes and at an initial cost of \$10, subject to further human review. A mere 1 percent reduction in the time required for a human to process a case would suffice to offset the costs associated with INACIA.
2.4. Vision for AI at the Court
TCU is an early adopter of AI technology, and INACIA was one of its major AI endeavors. INACIA will thus shape future AI adoption at the court. The current expectation is that AI will change how TCU handles cases from a sequential, step-by-step approach to a more holistic approach in which filings are evaluated simultaneously and connected to external resources necessary for decision-making from the outset. Effectively, this system would eliminate redundancies and streamline case processing without requiring regulation change.
3. The INACIA Project
3.1. Overview
Figure 1 summarizes INACIA’s main case processing pipeline, showcasing the holistic nature of case analysis. While there is a direct, sequential flow (i.e., solid arrows) of tasks (i.e., blue boxes) needed to produce case decision documents, the dashed lines show the association (or dependency) between tasks. The lighter grey boxes indicate data objects needed to start processing or to be produced as intermediate or final products for other TCU systems. In the next subsections, we describe each element of this pipeline in detail.
Figure 1. INACIA pipeline.
3.2. Inputs
The pipeline starts off ingesting all documents available at case submission (“Case documents” grey box). There are two sets of documents in the submission package: (1) the main document is the representação (in Portuguese), which is in nature similar to a civil complaint. It is a formal document laying out the claims of wrongdoing and why TCU should take the case; and (2) supporting documents containing the evidence of wrongdoing as presented by the plaintiff (e.g., if the plaintiff alleges that the defendant has not complied with a contract, the contract is a document that supports the case).
Both the representação and supporting documents are stored in multiple formats, and they can be classified in one of two categories: (1) searchable files:electronically-created files with searchable content (e.g., native PDF document or Word document printed as PDF); (2) unsearchable files:scanned physical files or images whose content is not searchable (e.g., a photo of a contract or signature). Since INACIA uses the textual content of case documents to perform its tasks, we needed to implement two data extraction solutions to handle all files stored by TCU. These solutions needed to deliver high-quality extraction at a low cost for the solution to be sustainable once TCU takes the system to production.
Most files in TCU’s case database are searchable and primarily stored as PDF documents. Thus, the first implemented solution is a standard PDF extraction pipeline using the Apache Tika framework (Apache, [n. d.]), which can handle all types of searchable files in TCU’s case database. Besides being an open-source tool (thus meeting the low-cost criterion), Tika delivered better results in text extraction than PyPDF2 (Fenniak, [n. d.]), another popular framework for processing searchable PDF documents. We compared and contrasted Tika and PyPDF2 in the following ways: (1) which framework was better at extracting textual content and its structure (e.g., a two-column document processed as separate, two columns instead of a single column with left-to-right text mixing up the two columns). (2) which framework was better at extracting characters from different encoding formats (e.g., those not present in UTF-8, Latin-1, or Windows-1252 encodings).
Tika performed better in both criteria – we particularly liked that it reports the count of invalid characters in the processed text as extraction metadata, allowing us to create a measure of extraction quality by easily computing the percentage of invalid characters over all characters extracted.
Whenever Tika fails in text extraction (e.g., it reported zero searchable characters after processing the text), INACIA then deploys the second text extraction solution, which relies on Microsoft’s Azure Form Recognizer service for Optical Character Recognition (OCR) (Microsoft, [n. d.]). Form Recognizer will ingest any unsearchable file and return the extracted text in plain format. We also compared Form Recognizer to other industry-leading OCR frameworks (Tesseract (Smith, 2007), and Google Cloud Vision (Google, [n. d.])) using the high-quality and low-cost criteria. While they are similarly priced, Form Recognizer had higher-quality results and richer extraction metadata (e.g., information on the position and content of tables).444 To assess text extraction from either document type, we evaluated results qualitatively. We manually compared extracted elements from case documents, such as the position of tables, columns, and text, allowing for a more nuanced understanding of the strengths and weaknesses of each tool in practical scenarios. We evaluated a sample of documents from TCU’s database containing all cases and their documents between 2019 and 2022. We pulled out a minimum of five documents per year. Within these documents, we had at least one document of easy, medium, and hard extraction, determined by the number of pages and their content as follows: (1) Easy: up to 10-page documents with minimal tables; (2) Medium: up to 25-page documents with structured data or images; (3) Hard: over 25-page documents with structured and unstructured data elements, images, or handwritten content.
3.3. Extraction of Basic Information and Allegations and Requests
The first two tasks (blue boxes in Figure 1) are the extraction of case summary information (“Basic information extraction” blue box), the allegations of wrongdoing made by the plaintiff (allegations), and what action they ask TCU to take (requests) (“Extraction of allegations and requests” blue box). We describe each task in the sections below.
3.3.1. Basic Information Extraction
Figure 2. Approach used for Basic Information Extraction.
Figure 2 illustrates the basic information extraction process. It is a structured and systematic process relying on prompt engineering and retrieval-augmented generation (RAG) techniques (Lewis et al., 2020): we will first feed the LLM a structured list of the information we need to extract from case documents (“Pre-defined form” green box). This list was constructed using older cases and their representação containing the information TCU needs to retain for record-keeping purposes 555 “Basic information” examples are the case ID, case type, plaintiff name, plaintiff ID, etc. . We also send over detailed explanations of what each of the variables is. Finally, we append to our combined prompt from above the first two pages of a case’s representação (“Main document” gray box), which should contain most of the information needed.
We deploy our RAG workflow when we cannot find all necessary information in the first pages of the representação. Along with extracting all textual content from case documents, we created searchable indexes of case documents. We allowed the LLM to reason through the form and query such indexes if it did not find the information in the first two pages of the main case document. An example that helps illustrate this mechanism is that TCU staff often need to look at federal agencies’ procurement documents to evaluate whether they have followed all necessary rules when contracting external vendors. INACIA can search supporting documents of a case to find procurement contracts and retrieve the necessary information (the rightmost side of Figure 2 illustrates this process). We use standard industry-leading search algorithms, such as BM25 (Robertson et al., 1994), to retrieve the relevant passages from the indexed case documents. The LLM then processes these passages to extract answers, ensuring a streamlined and consistent approach. The final step is retrieving the remaining information using BM25 and manually crafted queries. The resulting information is then fed into the LLM for the final extraction, completing the form. In this workflow, the LLM functions not as an intelligent agent but rather as a tool within the pipeline for interpreting search results.
3.3.2. Extraction of Allegations and Requests
An important step in our pipeline is to assess whether a case’s claims and supporting evidence are strong. Often, TCU staff can turn down the review of a case when there is weak evidence of wrongdoing when the plaintiff cannot straightforwardly articulate their claims that the defendant has broken the law in some shape or form, or even when the plaintiff does not clearly states their request for TCU action. As with other tasks, extracting allegations and actions requested by plaintiffs is a time-consuming task suited for LLM processing.
Figure 3. Prompt used to extract the allegations and requests.
## Instructions
You are an intelligent agent capable of reasoning and interpreting legal documents. You are given a case main document and an example of the allegations and requests to be extracted from the case main document. You must extract the allegations and requests from the case’s main document.
## Example
You must provide a list of allegations and requests similar to the example provided below:
[EXAMPLE TEXT GOES HERE.]
## Rules
You must follow the rules below:
1. Identify the allegations and requests presented by the plaintiff. They can be related to various aspects, including, but not limited to, the violation of a law, regulation, or contract.
2. Enumerate the allegations and requests in the same order as they appear in the case main document.
[CASE MAIN DOCUMENT TEXT GOES HERE.]
Similar to the basic information extraction task, we also implemented a combination of prompt engineering plus RAG to extract the relevant information (see Figure 3). We provide the model with an example of the allegations and requests to be extracted. The model extracts the allegations and requests from the case’s representação. We used OpenAI’s GPT-4 model with a context length of 32,000 tokens. The main advantage of this model is that it does not require fine-tuning or any additional training. The prompt contains instructions, an example of the allegations and requests to be extracted, a set of rules the LLM must follow, and the relevant case documents.
The data extracted in this step (allegations and requests) are used in three ways: (1) for the admissibility examination, where INACIA reviews cases for elements that would make such case “of public interest” and whether the claims are strong enough to hear the case; (2) for drafting the final decision document (called “initial instruction” or instrução inicial in Portuguese and cf. Section 3.7); (3) for the Fumus Boni Iuris analysis, which means evaluating a case for the likelihood of success on its merits (cf. Section 3.6).
3.4. Admissibility Examination
The next step in the INACIA project is to conduct an admissibility examination (“Admissibility exam” blue box in Figure 1). The admissibility examination is a preliminary examination conducted by auditors to determine whether the claims meet the minimum legal criteria to proceed. The TCU defines five criteria that must be met for a claim to be admissible:
  1. (1)
    Legitimacy: the case should be submitted by a legitimate plaintiff (as defined in TCU’s bylaws);
  2. (2)
    Competency: whether the wrongdoing falls within the jurisdiction of TCU (“competency” is the literal translation from Portuguese);
  3. (3)
    Existence of evidence: whether there is enough evidence supporting plaintiff’s claims;
  4. (4)
    Existence of public interest: whether the case is of public interest (as determined by potential (non-)monetary damage to the federal government).
  5. (5)
    Clear writing and language: whether the representação has been written clearly and objectively.
As shown in Figure 4, we used an approach similar to the one used for basic information extraction. We applied an LLM as an intelligent agent capable of reasoning and interpreting legal documents. However, instead of feeding the LLM the first two pages of the case’s representação, we allow the model to search for information in case documents using a search engine – including four external databases: (1) TCU’s jurisprudence database; (2) TCU’s statutes and Federal law; (3) TCU’s internal codes and other documents; and (4) the case’s documents.
Figure 4. Approach used for Admissibility Examination.
The admissibility examination is initiated by feeding the LLM one admissibility criterion. Then, the LLM searches for information in case documents and external databases to determine whether the case meets such admissibility criterion. For example, consider the admissibility criterion of legitimacy. The agent may search for the plaintiff’s information in the case documents and for information about who can file a claim in the external databases (e.g., the court’s internal codes). If the agent finds information that confirms a plaintiff’s legitimacy, the agent concludes that the claim meets the admissibility criterion. Otherwise, the agent concludes that the claim does not meet the admissibility criterion. This process is repeated for each admissibility criterion.
In this task, the agent can compose queries to search for information in case documents and external databases. For example, the agent may want to compose the query quem tem legitimidade para apresentar uma representação ao TCU? (i.e., “who has legitimacy to file a claim to the TCU?”) to search for information about who can file a claim in external databases. If that’s the case, the agent should not be limited to pre-defined queries and keyword-based queries. Therefore, the search engine, such as our implemented solution combining the BM25 algorithm and the monoT5666The monoT5 re-ranker is a neural network that re-ranks the documents based on the query and the document’s content. The re-ranker is a T5 model (Raffel et al., 2020) fine-tuned to indicate whether a document is relevant to a given query. The model produces a score between 0 and 1, indicating the document’s relevance to the query. This score is used to re-rank the documents. re-ranker (Nogueira et al., 2020). The BM25 algorithm is used to retrieve the documents that contain the words in the query. The monoT5 re-ranker is then used to re-rank the top 1,000 documents retrieved by the BM25 algorithm.
In addition to searching, we used an approach based on the Visconde system (Pereira et al., 2023) to reason through the information retrieved from case documents and external databases. Visconde is a multi-document question-answering system integrating information retrieval and LLMs to answer questions. The reasoning step is based on the Chain-of-Thought approach (Kojima et al., 2022) and produces a paragraph explaining how the documents support the model’s answer.
3.5. Periculum in mora Analysis
The Periculum in Mora (Latin for danger in delay) analysis is an essential part of INACIA, as it relates to the urgency with which TCU needs to take action in any particular case. It determines whether immediate action is required to prevent further damage or prejudice to the parties involved.
The analysis, however, is relatively straightforward. We carry out a simple search for exact matches of minuta de contrato (draft contract) in all case documents after a case has been assigned to a TCU staff, except when the case contains fewer than five documents – in which case all documents are considered. This search is essential to flag documents mentioning the draft contract so they are not included in subsequent steps. Two key rules are used to standardize the decision on whether there is Periculum in mora: (1) the existence of an active or signed contract – there is no danger in delay since TCU cannot prevent a contract from coming into force; (2) and (2) the existence of events that prove delay, cancellation, or impediment in contracting – there is no danger in delay because the execution of a contract has been halted. The verification of these rules occurs step-wise. We first verify whether there is any indication that a goods or service provision contract has been signed or is active. The verification is done directly from a keyword search and indirectly by asking the LLM to reason through the documents and verify the existence and status of procurement contracts. The second step is verifying any event delaying, canceling, or suspending a contract from entering into force. Here, we do not perform a direct keyword search for such information; we send documents and prompts straight to the LLM. The outcome of these searches provides an answer regarding the existence of an active contract and any delays or impediments. Finally, the findings are synthesized into a dictionary containing information about the presence of an active contract or delays, along with the document IDs from each search, and a decision on whether there is Periculum in mora, which can be either “rejected” or “accepted”.
It is important to note that the role of the LLM is limited to receiving search results and determining the existence of an active contract or impediments and delays. The LLM does not engage in function calling or actions like the ReAct framework (Yao et al., 2023), where the model defines actions to be taken. The ReAct framework is a paradigm that leverages reasoning and acting within language models to tackle a variety of language, reasoning, and decision-making tasks. The decision-making flows are hard-coded and not subject to the LLM’s discretion. After the decision is made, the result is passed onto the LLM along with a prompt laying out decision rules, enabling the LLM to draft the text that will be included in the final decision document (the “instruction”). This process ensures that the conclusions drawn from the Periculum in mora analysis are accurately reflected in the official documentation, preserving the integrity of TCU’s case review process.
3.6. Fumus boni iuris Analysis
The Fumus boni iuris analysis (Latin for smoke of good law) is also an essential part of INACIA; it refers to the “appearance of good law” or the likelihood of success on the merits of the case. This principle is applied to assess whether the plaintiff’s claims appear to be grounded in federal law and to have a reasonable chance of success should the case go to trial. INACIA uses its LLM to analyze the substance of the allegations and the legal principles they invoke. This involves thoroughly examining the legal arguments, jurisprudence, and relevant statutory laws. The LLM assesses the coherence of the allegations with existing legal norms and precedents, providing an initial legal opinion on the case’s merit.
For the Fumus boni iuris analysis, we used an approach similar to the one used for the admissibility examination. We used the LLM as an intelligent agent capable of reasoning and interpreting legal documents. However, instead of feeding the LLM with one of the admissibility criteria, we feed the LLM with the allegations extracted in the pipeline task “allegations and requests”. The LLM then searches for information in case documents and external databases to determine whether there is Fumus boni iuris.
The process for the Fumus boni iuris analysis is divided into two steps. In the first step, we feed the agent the allegations extracted in the previous project stage and ask him to classify each allegation as grounded in law, not grounded in law, or inconclusive. We use an approach based on the ReAct framework (Yao et al., 2023). We use the reasoning and acting components of the ReAct framework to instruct the model to classify the allegations. The reasoning component is used to reason through the allegations and plan the actions to be taken. The acting component executes the actions planned by the reasoning component, which includes searching the case documents and the external databases. One of the external databases used in this project stage is TCU’s selection of statutes and Federal law. The model searches for information about the laws invoked by the allegations in this database. The agent also searches for information about the allegations in the TCU’s jurisprudence database. The model then classifies the allegations as grounded in law, not grounded in law, or inconclusive.
3.7. Recommendations
The INACIA pipeline concludes with recommendations to the adjudication panel drafted by TCU staff assigned to review each. We call this the “Recommendations” stage. This pipeline task is relatively straightforward and recasts the other outputs from each task to construct a cohesive and coherent decision document to be sent over to the panel of judges who will issue the final ruling on a case. To write case recommendations, INACIA uses the following components:
  • The output from the admissibility examination, i.e., whether the case is admissible and the explanation of such decision.
  • The outcome of the periculum in mora and fumus boni iuris analyses, i.e., whether there is any danger in delay and the case is likely to be successful based on its merits.
  • The analysis of the case’s merits, with the details on claims, their relationship with the jurisprudence and statutory law, and recommended courses of action.
Each of these inputs is then processed by the LLM, which had been previously instructed with guidelines and examples of recommendations authored by experienced auditors. The LLM generates paragraphs for each recommendations section, ensuring the final text is coherent and accurately reflects the decisions and analyses. The end product is a structured document that is broken down into several sections, each on different aspects of the case, such as the case information, decisions on admissibility, merits, etc.
This modular construction of the final decision document makes sense both from the legal and the computational sense. From the legal perspective, it allows human reviewers to parse out the document into smaller chunks and use/correct them as they see fit. From the computational side, it means we do not have to reprocess everything in later stages if we need a small correction in one of the earlier stages of INACIA. This approach provides efficiency, especially in cases where, despite a potential flaw in one section, the rest of the information remains valid and useful.
The Recommendations task is thus a critical element of INACIA, marking the transition from analysis to actionable insights and ensuring that the outcomes of the case review process are communicated with clarity and precision, consistent with the standards upheld by the TCU.
4. Experiments
This section describes the experiments we conducted to evaluate the INACIA pipeline. Although INACIA has a variety of project tasks, we only evaluated the recommendations task, which represents the system’s final and most critical output. We first describe the dataset we used in the evaluation and then describe the evaluation techniques we used. Finally, we describe the results of the experiments.
4.1. Dataset Creation and Curation
A validation dataset was created to assess the quality of the analyses performed by INACIA. This dataset consists of a fully annotated table with the expected results for each pipeline step. In this study, we only worked with closed cases. This means the document of interest for evaluation (the decision document called instruction) was already available for all cases. This document is a searchable PDF written by TCU staff with expertise on the case subject. It is written in natural language and divided into five main sections: basic information, claims, and requests, admissibility examination, precautionary measure analysis (periculum in mora and fumus boni iuris, and recommendations. We used instructions from closed cases to build the validation dataset.
We extracted various elements from the instruction text to build the validation database. Each element followed its extraction strategy, as the information had different data structures. For basic information, GPT-4 was used to extract the data, such as names of parties, the amount involved, contract duration, and case summary, among others. In total, 26 basic variables were extracted. Before using GPT-4 in all cases, a random sample of 30 cases was manually analyzed to check if the data was being extracted with quality. This analysis resulted in 100% accuracy for the main information of interest.
For the claims and requests piece, the text was directly extracted from the instruction document, using regular expressions to isolate the section of interest. As the instruction documents considered are standardized, this extraction presented no errors. The same strategy was applied to the recommendations.
In the case of the admissibility examination, there are five criteria of interest, as described in Section 3.4. For each criterion, the labels “Yes (met)”, “No (not met)”, “Partial (partially met)”, or “Not applicable” were extracted using regular expressions. This information is systematically annotated in the instructions; hence, no extraction errors or manual corrections were needed.
The same technique of the admissibility examination was applied to the analysis of granting precautionary measures. There are three main criteria of interest: periculum in mora (danger in delay), reverse periculum in mora, and fumus boni iuris (legal plausibility). In all these cases, the possible outcomes of the analysis are “Dismissed (rejected)”, “Configured (accepted)”, “Inconclusive”, or “Not analyzed” (for example, when the case is not known). These data were extracted using regular expressions and did not present extraction errors.
Finally, a validation table was produced for all the cases of interest, containing basic information, claims, and requests, admissibility examination, analysis for granting precautionary measures and forwarding proposals. The final validation database effectively considered was a sample of 122 cases from all the over 1000 cases analyzed. The selection of cases that would be part of the final validation table was based on both some additional scope filters, focusing on cases with simpler analysis, and a random sample of the cases resulting from the filters.
Although we collect and validate a wide range of information, including basic case details, claims, admissibility assessments, and so on, not all results carry equal significance. We emphasize the evaluation of the recommendations, as it is the last and most important step of the instruction process.
4.2. Evaluation
For evaluating the INACIA main output, we used the validation dataset described in the previous section. The evaluation aims to compare the outcomes produced by the INACIA pipeline with the expected results previously produced by human auditors. This comparison could be done using term-matching metrics for these tasks, such as BLEU, Rouge, etc. However, these metrics are unsuitable for these tasks because the text generated by INACIA may not be the same as the text in the reference form. Instead, we used the LLM to evaluate if the candidate text is consistent with the reference text. Instead of asking the model to classify the candidate text as consistent or inconsistent with the reference text, we employ a more comprehensive approach.
Evaluating text generated by LLMs is a challenging task. While metrics such as BLEU and ROUGE are commonly used to evaluate text generation tasks, they have shown low correlation with human judgment (Fabbri et al., 2021). Approaches like GPTScore (Fu et al., 2023), and G-eval (Liu et al., 2023) have been proposed to address this issue by using LLMs to evaluate the generated text. They have been shown to correlate more with human judgment than traditional metrics when evaluating text summarization. However, these approaches are unsuitable for evaluating the text generated by the INACIA pipeline because the generated text is not a summary of a document. Instead, the generated text is the auditor’s recommendations.
For example, the objective of evaluating the recommendation generation is to assess whether the recommendations produced by INACIA are consistent with those produced by a human auditor. Intuitively, we want to evaluate if the candidate text (INACIA) contains the same elements as the reference text (auditor). For this, we prompted OpenAI’s GPT-4 with the reference recommendations from the validation dataset, asking it to compose a checklist of the elements that should be present in the recommendations. We then prompted the model with the candidate recommendations and asked it to evaluate if they contained the checklist elements. Then, we can count the number of checklist elements present in the candidate recommendations and use this number to evaluate the quality of the extracted information. For example, if the checklist contains ten elements and the candidate recommendations contain eight elements, we can say that the candidate recommendations are 80% consistent with the reference recommendations.
We also used the candidate document to prompt the model to compose a checklist of the elements that should be present in the candidate document. We then prompted the model with the reference document and asked it to evaluate if it contained the checklist elements. Then, we can count the number of checklist elements present in the reference document and use this number to evaluate the quality of the extracted information. This way, we have two scores: the precision, which indicates how consistent the candidate document is with the reference document, and the recall, which indicates how consistent the reference document is with the candidate document. We then use the F1 score to evaluate the quality of the extracted information. The F1 score is the harmonic mean of the precision and recall. We used the same approach to evaluate the recommendation generation.
This approach has shown to have a good performance in assessing the quality of the extracted text in the TCU context. To evaluate it, we asked an auditor to analyze five examples in which the score is too low (less than 0.3) and too high (more than 0.7). The auditor concluded that the examples with low scores were not consistent with the reference text, and those with high scores were consistent with the reference text. This shows that the approach is suitable for evaluating the quality of the extracted text in the TCU context. Yet, we evaluated this approach on assessing the consistency of the automatically generated summary so we could compare it with GPTScore and G-eval. We used the SummEval Dataset (Fabbri et al., 2021) as a benchmark. The results shown in Table 1 demonstrate that our approach correlates more with human judgment than GPTScore and G-eval on assessing how consistent a summary is concerning its original text. It also has a higher correlation with human judgment, considering the overall quality of the summary. This shows that our approach is suitable for evaluating the quality of the automatically generated text in the TCU context.
Table 1. Check-eval results. Summary-level ($\rho$) Spearman and ($\tau$) Kendall-Tau correlations.
Method Consistency Relevance Coherence Fluency AVG
$\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$
G-eval 0.507 0.425 0.547 0.433 0.582 0.457 0.455 0.378 0.514 0.418
GPTScore 0.449 0.381 0.434 0.403 0.417
check-eval 0.605 0.570 0.5028 0.4202 0.563 0.461 0.490 0.446 0.623 0.493
5. Results
Table 2 presents the results of evaluating the recommendations generated by the INACIA pipeline using the checklist approach. This method assesses the alignment of the generated recommendations with the checklist of elements derived from the reference text. The table shows the mean, standard deviation (Std), minimum (Min), and maximum (Max) values for precision, recall, and the F1 score. These metrics provide insights into the consistency and completeness of the generated recommendations compared to the reference text. The average precision score of 0.592 indicates that, on average, 59.2% of the checklist elements were found in the generated recommendations. This suggests a moderate level of accuracy in capturing the necessary elements in the recommendations. The average recall score is 0.401, indicating that roughly 40.1% of the reference text elements were present in the generated recommendations. This lower recall suggests that the generated recommendations do not capture significant reference text elements. The average F1 score, the harmonic mean of precision and recall, is 0.429. This score reflects a balance between precision and recall but suggests room for improvement in both aspects. The standard deviation values for precision (0.212), recall (0.275), and F1 (0.225) show a certain degree of variability in the performance across different cases. Overall, these results suggest that while there are instances where the recommendations closely match the reference text, there is a noticeable variation in the quality of the generated recommendations. The relatively lower recall scores indicate a need for improvement in capturing all the necessary elements from the reference text in the generated recommendations.
Table 2. Results of the recommendations using the checklist approach.
Precision Recall F1
Mean 0.592 0.401 0.429
Std 0.212 0.275 0.225
Min 0.000 0.059 0.000
Max 1.000 1.000 0.933
6. Conclusions
In this paper, we presented INACIA, an innovative system designed to integrate Large Language Models (LLMs) into the Brazilian Federal Court of Accounts operational framework. Our experiments and analysis reveal that INACIA holds significant potential to transform and optimize the audit process by automating and streamlining case analysis and decision-making.
The utilization of LLMs for processing and analyzing case documents demonstrated a moderate level of accuracy in capturing essential elements and information from cases. The system’s ability to handle various stages of case processing, from basic information extraction to admissibility examination, precautionary measure analysis, legal plausibility assessment, and final recommendations generation, showcases its versatility and adaptability to the complex requirements of legal proceedings.
Moving forward, we will focus on enhancing INACIA’s capabilities through continuous learning and adaptation. Besides, we will explore implementing an auditing system based on INACIA’s architecture, allowing us to evaluate the system’s performance in a real-world setting. In conclusion, INACIA represents a significant step forward in applying artificial intelligence in the legal domain. Its development and implementation in the Brazilian audit courts can serve as a blueprint for similar initiatives globally.